OpenStudy (anonymous):

Find the values of x and y. 3x=4x-5

OpenStudy (nubeer):

hmm 1. subtract 3x on both sides..

OpenStudy (nubeer):

2. add 5 on both sides.

OpenStudy (anonymous):

Wait, if you subtract 3x on both sides on 3x=4x-5, the the 3x completely disappears so you're left with x-5.

OpenStudy (nubeer):

hmm to be precise u r left with 0=x-5.. right?

OpenStudy (anonymous):

Oh... My bad :) So, 5=x?

OpenStudy (nubeer):

yup.. there you go :)

OpenStudy (anonymous):

Soo.. On y+3=2y-1 .. the answer would be y=4?

OpenStudy (nubeer):

yes.

OpenStudy (anonymous):

Soo.. On 2x=y+3 with the y=4, the answer would be 3.5?

OpenStudy (nubeer):

yes

OpenStudy (anonymous):

Sweetness! :) Thanks.

OpenStudy (nubeer):

no problem at all ;)